# The Potential and Challenges of Evaluating Attitudes, Opinions, and Values in Large Language Models

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have sparked wide interest in validating and comprehending the human-like cognitive-behavioral traits LLMs may have. These cognitive-behavioral traits include typically Attitudes, Opinions, Values (AOV). However, measuring AOV embedded within LLMs remains opaque, and different evaluation methods may yield different results. This has led to a lack of clarity on how different studies are related to each other and how they can be interpreted. This paper aims to bridge this gap by providing an overview of recent works on the evaluation of AOV in LLMs. Moreover, we survey related approaches in different stages of the evaluation pipeline in these works. By doing so, we address the potential and challenges with respect to understanding the model, human-AI alignment, and downstream application in social sciences. Finally, we provide practical insights into evaluation methods, model enhancement, and interdisciplinary collaboration, thereby contributing to the evolving landscape of evaluating AOV in LLMs.Lead authors.
@@ 1 Introduction
Recent years have witnessed a remarkable improvement in the development and deployment of Large Language Models (LLMs), holding the promise of boosting various domains, from computer sciences to social sciences and beyond. Amid the excitement surrounding their capabilities lies an important question: How well do these LLMs capture and convey human cognitive-behavioral traits?
By drawing upon theories from social sciences (such as Katz, 1960; Rokeach, 1968; Ajzen, 1988; Bergman, 1998), we consider human cognitive-behavioral traits, in our case primarily Attitudes, Opinions, Values (AOV), the fundamental component of human cognition, shaping our perceptions, decisions, and interactions. By studying whether and how the LLM outputs reflect AOV and how these AOV compare to human AOV, we can better understand their potential to act as autonomous agents that could mirror human AOV. The AOV in LLMs also impact users in downstream applications, such as writing assistants Jakesch et al. (2023), and affect decision-making processes and perceptions Eigner and Händler (2024).
In recent studies, survey questionnaires that were originally used to estimate public opinions in the social sciences are now being popularly utilized to evaluate the opinions of LLMs and subsequently to study the alignment with human opinions (Santurkar et al., 2023; Hwang et al., 2023; Kim and Lee, 2024, inter alia). At the same time, the wide range of evaluation methods used to assess LLM responses has led to inconsistent outcomes, complicating reliable assessment of the models (Dominguez-Olmedo et al., 2023; Wang et al., 2024a, b, inter alia). However, this variability in evaluation methods has largely been overlooked—posing risks of missing subtleties in LLM performance, yielding incomplete or biased assessments. This oversight raises significant questions about the model’s true capabilities and its alignment with human opinions.
Motivated by the rising interest in studying the human-like traits of LLMs, in this paper, we present the first survey on the evaluation of AOV in LLMs. Before moving into the details, we first position our survey in the context of other relevant surveys and then show the framework of our survey.
@@
While there are no survey papers specifically on AOV in LLMs, some existing works have covered related questions. Simmons and Hare (2023) review works and provide a framework for using LLMs as comparative models for subpopulations to measure public opinions. Jansen et al. (2023) offer insights on employing LLMs in public opinion survey research, concluding that LLMs can enhance survey research. Vida et al. (2023) address the research gap in the ethical aspects of NLP surveying the literature on moral NLP, calling for a more rigorous discussion on the moral concept for NLP research. Hershcovich et al. (2022) provide a survey on NLP in cross-cultural contexts from a linguistic diversity angle, suggesting the need to preserve cultural values in models. There are also recent surveys on understanding “culture” in NLP Liu et al. (2024) and measuring “cultures” in LLMs Adilazuarda et al. (2024), both highlighting the future of culturally aware and adapted NLP techniques. These works have mainly explored topics like studying the cultural and moral aspects in NLP or improving public opinion research with LLMs. However, there has been a lack of dedicated studies focusing on AOV and especially on evaluating the AOV within LLMs.
@@
Since LLMs are pretrained on vast amounts of human data, it is reasonable to hypothesize that LLMs can reflect the AOV embedded in the data. But, for that to scale, we will need definitions of the terms AOV (WHAT is it? §2), then to summarize what has been explored on the AOV in LLMs so far (WHAT so far? §3), and know the pipeline used so far in research on how LLMs are queried for the AOV embedded within (HOW? §4). We then discuss the research directions (WHERE? §5) by highlighting the potential and challenges identified from existing works and the evaluation pipeline. In the end, we provide a call for action on what to do to make these approaches possible and reliable in the future (WHAT to do? §6).
@@ 2 Definitions
Next, we provide definitions for the three main concepts used in this paper: attitude, opinion, and value (WHAT is it?). According to Katz (1960), an attitude is a durable orientation toward some object, while an opinion is more of a visible expression of an attitude. For this paper, we examine the two concepts simultaneously following Bergman (1998), who considers the attitude and opinion as synonymous:
@@
3 An Overview of Related Works for AOV in LLMs
In this section, we present related recent works on the evaluation of AOV in LLMs (WHAT so far?). We categorize the works into two main groups: attitudes/opinions and values, reflecting the two dimensions of AOV we proposed. In addition, we include works with various topics that could also shine light on AOV in LLMs. A summary of the surveyed papers, along with the details on the paper selection process and an analysis of the model distribution can be found in the Appendix (§A.1, §A.2).
@@ 3.1 Attitudes/Opinions
The majority of recent work on evaluating opinions in LLMs is based on US-centric public opinion surveys. Argyle et al. (2023), Bisbee et al. (2023) and Sun et al. (2024) query the model with a prompt that encompasses the socio-demographics of real human participants using the American National Election Studies (ANES) surveys. Santurkar et al. (2023) use the American Trends Panel (ATP) survey from the Pew Research Center and create the dataset OpinionQA. The OpinionQA data set has also been used by Hwang et al. (2023) and Wang et al. (2024b). Similarly, Tjuatja et al. (2024) also use ATP data to study whether LLMs exhibit human-like response biases. There are various additional US-based surveys used to study LLMs’ AOV Dominguez-Olmedo et al. (2023); Kim and Lee (2024); Sanders et al. (2023); Lee et al. (2024a). Most of the papers found misalignment between LLM and human opinions and several observed left-leaning political bias in their comparisons.
@@
Although most work relies on the US context, a few studies focus on non-US countries or cross-national comparisons. von der Heyde et al. (2023) use data from German Longitudinal Election Study GLES (2019) and notice strong bias also in their use case (German election prediction). Kalinin (2023) uses the Survey of Russian Elites from 1993–2020 Zimmerman et al. (2023) and leverages LLMs to generate opinions like Russian elite individuals. Durmus et al. (2023) introduce the dataset GlobalOpinionQA based on questions and answers from cross-national surveys on diverse opinions on global issues across different countries and discover cultural and social biases of LLMs’ outputs.
Non-Public-Opinion Polls.
@@ 3.2 Values
Moral values are an important component of value orientation. Notably, the Moral Foundations Theory [^1] The Moral Foundations Theory Graham et al. (2011) identifies five foundations (Care, Fairness, Loyalty, Authority, Purity) to explain shared moral themes across populations Abdulhai et al. (2023). The Moral Foundations Questionnaire Graham et al. (2011) scores these five foundations. Graham et al. (2018) has been applied in several studies to assess models’ moral values. For instance, Simmons (2023) investigates moral biases in LLMs using this theory, demonstrating that these models exhibit moral biases when prompted with certain political identities. Haemmerl et al. (2023) explore moral foundations in multilingual LLMs. Abdulhai et al. (2023) show that LLMs can be properly prompted to exhibit specific moral foundations. Findings on the evaluation of morals in LLMs are sometimes inconsistent. While Talat et al. (2022) claim that the models exhibit fluctuating moral values, Fraser et al. (2022) find that the models’ moral values align consistently with their training data. Additionally, Bonagiri et al. (2024) note that some LLMs exhibit inconsistency when responding to moral questions phrased differently but in semantically equivalent ways.
@@
There are a few curated evaluation datasets for values in LLMs, such as ETHICS Hendrycks et al. (2023), MoralChoice Scherrer et al. (2024), MoralExceptQA Jin et al. (2022), ValuePrism Sorensen et al. (2024), Moral Consistency Corpus Bonagiri et al. (2024). A few frameworks have been established to assess the ethical reasoning capability of LLMs, such as SocialChemistry101 Forbes et al. (2020), Delphi Jiang et al. (2022a), the Framework for ‘in-context’ Ethical Policies Rao et al. (2023), Moral Graph Elicitation Klingefjord et al. (2024), SaGE to measure moral consistency Bonagiri et al. (2024), as well as moral dilemmas and value statements Rao et al. (2023); Agarwal et al. (2024). Ren et al. (2024) provide an evaluation pipeline ValueBench to probe value orientations encompassing 453 value dimensions. These resources and frameworks collectively enhance our ability to evaluate and understand the values embedded in LLMs.
@@ 3.3 Other Related Topics
In addition to the two main categories, several studies investigate related topics that indirectly also reveal the AOV reflected in LLMs. These include: i) trustfulness, which is closely related to AOV as it reflects the model’s alignment to human values on truth and honesty Lin et al. (2022); Joshi et al. (2024), ii) theory-of-mind, which explores the ability of LLMs to understand and predict human thoughts and emotions Sap et al. (2022); Li et al. (2023b); Kosinski (2024), iii) persona and personality, of which findings highlight the models’ ability to reflect human-like attitudes and values through their generated personas Miotto et al. (2022); Kovač et al. (2023); Caron and Srivastava (2023); Cheng et al. (2023a, b); Jiang et al. (2024); Shu et al. (2024), iv) sentiment Deshpande et al. (2023); Beck et al. (2024b); Hu and Collier (2024), and v) mixed topics spanning politics, philosophy and personality (e.g. Perez et al., 2023).
@@
4 How LLMs Are Queried for AOV
After defining the core concepts and discussing related works, we now provide details of the pipeline on how LLMs were queried for AOV so far (HOW?) to motivate our later discussion on gaps. Based on the surveyed works, we categorize the evaluation process in a taxonomy into four main stages: i) input, ii) model, iii) output, and iv) evaluation, as illustrated in Figure 1.
@@ 4.1 Input
In this section, we show methods for formatting input data before feeding them into the model. Several examples of the task design for the input can be found in the Appendix §A.3.
@@
To test the robustness and consistency of the model’s outputs, perturbations have been applied to the input to test the human-like response biases of the model. The most common way is to perturbate the order of the choices in close-ended questions Lu et al. (2022); Hartmann et al. (2023); Kovač et al. (2023); Dominguez-Olmedo et al. (2023); Tjuatja et al. (2024); Wang et al. (2024b); Shu et al. (2024), or to paraphrase the original questions Feng et al. (2023); Hartmann et al. (2023); Shu et al. (2024); Röttger et al. (2024); Bonagiri et al. (2024); Wright et al. (2024). Tjuatja et al. (2024) propose response bias modifications (e.g. order swapping) and non-bias perturbations (e.g. letter swapping and typos), which are also employed in Wang et al. (2024a). In addition, modifying instruction prompt wording is another perturbation approach. Cao et al. (2023) change questions from the second to the third person, while Kovač et al. (2023) and Ceron et al. (2024) prepend a system message in the second person to the question. Hwang et al. (2023) add a Chain-of-Thought (CoT, Wei et al., 2023) style prompt wording to the questions. Röttger et al. (2024) apply prompts that force LLMs to choose a multiple-choice answer.
@@ 4.2 Model
Fine-Tuning and Inference.
Some studies utilize the fine-tuning approach to align LLMs with specific viewpoints by training them on data containing those opinions (e.g. partisan Twitter data, parliamentary debates), and during the inference period then evaluate these fine-tuned models on test sets (e.g. questionnaires for human public opinion polls) Jiang et al. (2022b, a); Joshi et al. (2024); Chalkidis and Brandl (2024); Kim and Lee (2024). These works showed that the fine-tuned models can represent the opinions behind the training data.
@@
In multi-turn inference, the process is usually chain-wise or conducted by multiple agents. Perez et al. (2023) instruct LLMs to write yes/no questions with multiple stages of LM-based generation and filtering. Several works Jin et al. (2022); Jiang et al. (2023b); Yang et al. (2023) incorporate CoT processes to complete questionnaires in a multi-turn dialogue manner, while Baltaji et al. (2024) use multi-agent LLM systems for inter-cultural collaboration and debate, analyzing opinion diversity before and after agent discussions, based on previous research on social behaviors in LLM agents Li et al. (2023b); Zhang et al. (2024).
@@ 4.3 Output
After defining the inputs and models and feeding the input into the model, we can now address the output side. There are two main ways for output extraction: logits-based output and text-based output.
@@
The first token logits of LLM outputs have been commonly used in multiple-choice question settings to transform the open-ended nature of LLM outputs into expected options, as in Hendrycks et al. (2021b); Santurkar et al. (2023); Dominguez-Olmedo et al. (2023); Chalkidis and Brandl (2024); Kalinin (2023); Beck et al. (2024b); Lin et al. (2022). This method involves calculating the log probabilities for answer options (e.g. ‘A’, ‘B’, ‘C’). The option with the highest log probability is then selected as answer.

### v3

@@ Abstract
Recent advances in Large Language Models (LLMs) have sparked wide interest in validating and comprehending the human-like cognitive-behavioral traits LLMs may capture and convey. These cognitive-behavioral traits include typically Attitudes, Opinions, Values (AOVs). However, measuring AOVs embedded within LLMs remains opaque, and different evaluation methods may yield different results. This has led to a lack of clarity on how different studies are related to each other and how they can be interpreted. This paper aims to bridge this gap by providing a comprehensive overview of recent works on the evaluation of AOVs in LLMs. Moreover, we survey related approaches in different stages of the evaluation pipeline in these works. By doing so, we address the potential and challenges with respect to understanding the model, human-AI alignment, and downstream application in social sciences. Finally, we provide practical insights into evaluation methods, model enhancement, and interdisciplinary collaboration, thereby contributing to the evolving landscape of evaluating AOVs in LLMs.Lead authors.
@@ 1 Introduction
Recent years have witnessed a remarkable improvement in the development of Large Language Models (LLMs), holding the promise of boosting various domains, from computer sciences to social sciences and beyond Ziems et al. (2024). Amid the excitement surrounding their capabilities, when we take a human-centric perspective, an important question lies: How well do these LLMs capture and convey human cognitive-behavioral traits?
By drawing upon traditional theories from the social sciences (such as Katz, 1960; Rokeach, 1968; Ajzen, 1988; Bergman, 1998), we consider human cognitive-behavioral traits, in our case specifically Attitudes, Opinions, Values (AOVs), as fundamental components of human cognition, shaping our perceptions, decisions, and interactions. By examining whether and how LLM outputs reflect AOVs, and comparing these AOVs to those of humans, we can gain deeper insights into the models’ capacity to function as autonomous agents mirroring human AOVs. The AOVs in LLMs also impact users in downstream applications, such as writing assistants Jakesch et al. (2023), and affect decision-making processes and perceptions Eigner and Händler (2024).
In recent studies, survey questionnaires that were originally used to estimate public opinions in the social sciences are now being popularly utilized to evaluate the opinions of LLMs and subsequently to study the alignment with human opinions (Santurkar et al., 2023; Hwang et al., 2023; Kim and Lee, 2024, inter alia). At the same time, the wide range of evaluation methods used to assess LLM responses has led to inconsistent outcomes, complicating reliable assessment of the models. For instance, Röttger et al. (2024) demonstrate that different prompting methods lead to different results; Wang et al. (2024b) show that output extraction methods can yield misaligned results. However, this variability in evaluation methods has sometimes been overlooked in real usecases—posing risks of missing subtleties in LLM performance, yielding incomplete or biased assessments. This oversight raises questions about the model’s true capabilities and its alignment with human opinions.
Motivated by the rising interest in studying the human-like traits of LLMs, in this paper, we present the first survey on the evaluation of AOVs in LLMs. Before moving into the details, we first position our survey in the context of other relevant surveys and then show the framework of our survey.
@@
While there are no survey papers specifically on AOVs in LLMs, some existing works have covered related questions. On opinions, Simmons and Hare (2023) review frameworks for using LLMs to model subpopulations and measure public opinions. Similarly, Jansen et al. (2023) explore the use of LLMs in public opinion research, noting their potential to enhance survey methods. On values, Vida et al. (2023) survey moral NLP, emphasizing the need for more rigorous discussions on ethical concepts. Hershcovich et al. (2022) examine NLP in cross-cultural contexts, urging the preservation of cultural values. Additionally, recent surveys on understanding “culture” in NLP Liu et al. (2024) and measuring “cultures” in LLMs Adilazuarda et al. (2024) highlight the importance of culturally aware techniques. These works have mainly explored topics like improving public opinion research with LLMs or studying the cultural and moral aspects in NLP. However, there has been a lack of dedicated studies focusing on AOVs and especially on evaluating the AOVs within LLMs.
@@
Since LLMs are pretrained on vast amounts of human data, it is reasonable to hypothesize that LLMs can reflect the AOVs embedded in the data Durmus et al. (2024). But, for that to scale, we will need definitions of the terms AOVs (WHAT are they? §2), then to summarize what has been explored on the AOVs in LLMs so far (WHAT so far? §3), and know the pipeline used so far in research on how LLMs are queried for the AOVs embedded within (HOW? §4). We then discuss the research directions (WHERE? §5) by highlighting the potential and challenges identified from existing works and the evaluation pipeline. In the end, we provide a call for action on what to do to make these approaches possible and reliable in the future (WHAT to do? §6).
@@ 2 Definitions
Next, we provide definitions for the three main concepts used in this paper: attitude, opinion, and value (WHAT are they?). According to Katz (1960), an attitude is a durable orientation toward some object, while an opinion is more of a visible expression of an attitude. For this paper, we examine the two concepts simultaneously following Bergman (1998), who considers the attitude and opinion as synonymous:
@@
3 An Overview of Related Works for AOVs in LLMs
In this section, we present related recent works on the evaluation of AOVs in LLMs (WHAT so far?). We categorize the works into two main groups: attitudes/opinions and values, reflecting the two dimensions of AOVs we proposed. In addition, we include works with various topics that could also shine light on AOVs in LLMs. A summary of the surveyed papers, along with the details on the paper selection and an analysis of the model distribution can be found in the Appendix (§A.1, §A.2).
@@ 3.1 Attitudes/Opinions
The majority of recent work on evaluating opinions in LLMs is based on US-centric public opinion surveys. Argyle et al. (2023), Bisbee et al. (2023) and Sun et al. (2024) query the model with a prompt that encompasses the socio-demographics of real human participants using the American National Election Studies (ANES) surveys. Santurkar et al. (2023) use the American Trends Panel (ATP) survey from the Pew Research Center and create the dataset OpinionQA. The OpinionQA data set has also been used by Hwang et al. (2023) and Wang et al. (2024b). Similarly, Tjuatja et al. (2024) also use ATP data to study whether LLMs exhibit human-like response biases. There are various additional US-based surveys used to study LLMs’ AOVs Dominguez-Olmedo et al. (2023); Kim and Lee (2024); Sanders et al. (2023); Lee et al. (2024a). Most of the papers found misalignment between LLM and human opinions and several observed left-leaning political bias in their comparisons.
@@
Although most work relies on the US context, a few studies focus on non-US countries or cross-national comparisons. von der Heyde et al. (2023) use data from German Longitudinal Election Study GLES (2019) and notice strong bias also in their use case (German election prediction). Kalinin (2023) uses the Survey of Russian Elites from 1993–2020 Zimmerman et al. (2023) and leverages LLMs to generate opinions like Russian elite individuals. Geng et al. (2024) conduct LLM experiments on European Social Survey (ESS). Durmus et al. (2024) introduce the dataset GlobalOpinionQA based on questions and answers from cross-national surveys on global issues across different countries and discover cultural and social biases of LLMs’ outputs.
Additional Data Sources Used.
@@ 3.2 Values
Moral values are an important component of value orientation. Notably, the Moral Foundations Theory [^1] The Moral Foundations Theory identifies five foundations (Care, Fairness, Loyalty, Authority, Purity) to explain shared moral themes across populations Abdulhai et al. (2023). The Moral Foundations Questionnaire Graham et al. (2011) scores these five foundations. Graham et al. (2018) has been applied in several studies to assess models’ moral values. For example, Simmons (2023) examine moral biases in LLMs, showing that these models exhibit bias when prompted with political identities. Haemmerl et al. (2023) explore moral foundations in multilingual LLMs, while Abdulhai et al. (2023) show that LLMs can be properly prompted to exhibit specific moral foundations. Findings on the evaluation of morals in LLMs are sometimes inconsistent: Talat et al. (2022) report fluctuating moral values, while Fraser et al. (2022) find alignment with training data. Bonagiri et al. (2024) highlight inconsistencies in responses to rephrased moral questions. Cross-culturally, Jinnai (2024) focus on aligning LLMs with Japanese commonsense morality, advocating for cross-cultural alignment.
@@
There are a few curated evaluation datasets for values in LLMs, such as ETHICS Hendrycks et al. (2023), MoralChoice Scherrer et al. (2023), MoralExceptQA Jin et al. (2022), ValuePrism Sorensen et al. (2024), Moral Consistency Corpus Bonagiri et al. (2024), and CMoralEval for Chinese LLMs Yu et al. (2024). A few frameworks have been established to assess the ethical reasoning capability of LLMs, such as SocialChemistry101 Forbes et al. (2020), Delphi Fraser et al. (2022), the Framework for ‘in-context’ Ethical Policies Rao et al. (2023a), Moral Graph Elicitation Klingefjord et al. (2024), SaGE to measure moral consistency Bonagiri et al. (2024), as well as moral dilemmas and value statements Rao et al. (2023a); Agarwal et al. (2024). Ren et al. (2024) provide an evaluation pipeline ValueBench to probe value orientations encompassing 453 value dimensions. These resources and frameworks collectively enhance our ability to evaluate and understand the values embedded in LLMs.
@@ 3.3 Other Related Topics
In addition to the two main categories, several studies investigate related topics that indirectly also reveal the AOVs reflected in LLMs. These include: i) trustfulness, which is closely related to AOVs as it reflects the model’s alignment to human values on truth and honesty Lin et al. (2022); Joshi et al. (2024), ii) theory-of-mind, which explores the ability of LLMs to understand and predict human thoughts Sap et al. (2022); Li et al. (2023b); Kosinski (2024); Strachan et al. (2024), iii) persona and personality, of which findings highlight the models’ ability to reflect human AOVs through their generated personas Miotto et al. (2022); Kovač et al. (2023); Rao et al. (2023b); Caron and Srivastava (2023); Cheng et al. (2023a, b); Jiang et al. (2024); Shu et al. (2024), iv) sentiment Deshpande et al. (2023); Beck et al. (2024b); Hu and Collier (2024), and v) mixed topics spanning politics, philosophy and personality (e.g. Perez et al., 2023).
@@
4 How LLMs Are Queried for AOVs
After defining the core concepts and discussing related works, we now provide details of the pipeline on how LLMs were queried for AOVs so far (HOW?) to motivate our later discussion on gaps. Based on the surveyed works, we categorize the evaluation process in a taxonomy into four main subcategories: i) input, ii) model, iii) output, and iv) evaluation, covering the four main stages of the evaluation process, as illustrated in Figure 1.
@@ 4.1 Input
In this section, we show methods for formatting input data before feeding them into the model. Several examples of the task design for the input can be found in the Appendix §A.3. Apart from the common direct input prompting, following two specific input formatting approaches have been applied.
@@
To test the robustness and consistency of the model’s outputs, perturbations have been applied to the input to test the human-like response biases of the model. The most common way is to perturbate the order of the choices in close-ended questions Lu et al. (2022); Hartmann et al. (2023); Kovač et al. (2023); Dominguez-Olmedo et al. (2023); Tjuatja et al. (2024); Wang et al. (2024b); Shu et al. (2024), or to paraphrase the original questions Feng et al. (2023); Hartmann et al. (2023); Shu et al. (2024); Röttger et al. (2024); Bonagiri et al. (2024); Wright et al. (2024). Tjuatja et al. (2024) propose response bias modifications (e.g. order swapping) and non-bias perturbations (e.g. letter swapping and typos), which are also employed in Wang et al. (2024a). In addition, modifying instruction prompt wording is another perturbation approach. Cao et al. (2023) change questions from the second to the third person, while Kovač et al. (2023) and Ceron et al. (2024) prepend a system message in the second person to the question. Hwang et al. (2023) add a Chain-of-Thought (CoT, Wei et al., 2022) style prompt wording to the questions. Röttger et al. (2024) apply prompts that force LLMs to choose a multiple-choice answer.
@@ 4.2 Model
Fine-Tuning and then Inference.
Some studies utilize the supervised fine-tuning approach (e.g. LoRA, Hu et al., 2022) to align LLMs with specific viewpoints by training them on data containing those opinions (e.g. partisan Twitter data, parliamentary debates), and during the inference period then evaluate these fine-tuned models on test sets of human data or curated benchmarks. Jiang et al. (2022b, a); Joshi et al. (2024); Chalkidis and Brandl (2024); Kim and Lee (2024); Jinnai (2024). These works showed that the fine-tuned models can represent the opinions behind the training data.
@@
In multi-turn inference, the process is usually chain-wise or conducted by multiple agents. Perez et al. (2023) instruct LLMs to write yes/no questions with multiple stages of generation and filtering. Several works Jin et al. (2022); Jiang et al. (2023b); Yang et al. (2023) incorporate CoT processes to complete questionnaires in a multi-turn dialogue manner, while Baltaji et al. (2024) use multi-agent LLM systems for inter-cultural collaboration and debate, analyzing opinion diversity before and after agent discussions, based on previous research on social behaviors in LLM agents Li et al. (2023b); Zhang et al. (2024).
@@ 4.3 Output
After defining the inputs and models and feeding the input into the model, we can now address the output side. There are two main ways for output extraction: logits-based and text-based output.
@@
The first token logits of LLM outputs have been commonly used in multiple-choice question settings to transform the open-ended nature of LLM outputs into expected options, as in Hendrycks et al. (2021); Santurkar et al. (2023); Dominguez-Olmedo et al. (2023); Chalkidis and Brandl (2024); Kalinin (2023); Beck et al. (2024b); Lin et al. (2022). This method involves calculating the log probabilities for answer options (e.g. ‘A’, ‘B’, ‘C’). The option with the highest log probability is then selected as answer.